# All-Electron Molecular Tunnel Ionization Based on the Weak-Field Asymptotic Theory in the Integral Representation

### v2

@@ IV.4 N2, HCOOH, and CH3F
Figure 6: (a) and (b): $|G_{00}|^{2}$ of CH3F obtained by OE-WFAT using HOMO(–) and HOMO(+), respectively. (c) and (d): $|G_{00}|^{2}$ of CH3F obtained by ME-WFAT using Dyson(–) and Dyson(+), respectively. (e) The shape of CH3F. (f) HOMO(–). (g) Dyson(–). (h) The values of $\mathcal{P}^{\prime}(i)$ of several MO with the largest contribution to Dyson(–). The above structure factors are based on the reference that $(\beta,\gamma)=(0^{\circ},0^{\circ})$ happens when the C–F bond coincides the $z$-axis.
@@
As the last example, we examine CH3F, a molecule having an equilibrium geometry that possesses a $C_{3v}$ symmetry, where the axis of the three-fold rotational symmetry coincides with the CF bond. The C–F and C–H bond lengths are $1.360$ and $1.080$, respectively, with the HCF angle of $108.93^{\circ}$. Its HOMO is doubly degenerate and can be identified based on their reflection symmetry with respect to the plane containing one of the three FCH planes, here, we denote the symmetric and antisymmetric functions as HOMO(+) and HOMO(–), respectively. We choose the reflection plane to be the $xz$-plane. Fig. 6(a) and 6(b) are the $|G_{00}|^{2}$ structure factors obtained by OE-WFAT using HOMO(–) and HOMO(+), respectively. The ground state of CH3F+ is also doubly degenerate, which is the consequence of the equal probabilities of removing an electron from either HOMO(+) or HOMO(–) of the neutral to form the ground state cation. This means that one may construct two “degenerate” Dyson orbitals, Dyson(+) and Dyson(–). The ME-WFAT $|G_{00}|^{2}$ structure factors using Dyson(–) and Dyson(+) are shown in Fig. 6(c) and 6(d), respectively. We see that the positions as well as the relative intensities are visibly different compared to OE-WFAT results (Fig. 6(a) and (b)), for example, the peaks are at smaller $\beta$ for ME-WFAT structure factors (Fig. 6(c) and 6(d)) and the middle bottom lobe in Fig. 6(d) (ME-WFAT) is at a lower intensity than the corresponding lobe in Fig. 6(b) (OE-WFAT). Again, we can understand this observation by looking at the shapes of the HOMO and Dyson orbital shown in Fig. 6(f) and (g), respectively, for the odd symmetry. The balancing of the upper and lower pairs of lobes of HOMO(–) (Fig. 6(f)) in the formation of the Dyson orbital (Fig. 6(g)) are due to the appreciable interference between HOMO(–) and HOMO-2(–) (HOMO-2 is also doubly degenerate) that has a comparable magnitude of $\mathcal{P}(i)$ (see Fig. 6(h)).
@@
As a final note regarding the Dyson orbital, the $\mathcal{P}(i)$ coefficients, which encode the contribution of the neutral molecule’s MO to the Dyson orbital, is independent of the orbital energies of these MO. This is worth mentioning because one might be tempted to think that $|\mathcal{P}(i)|$ should die down monotonically as a function of orbital energy difference relative to that of the orbital that dominates the calculated ionization channel (usually the HOMO for ground-state-to-ground-state channels). This is an incorrect view because these contributions only tell us how much of these MO is needed to build an optimal ionizing orbital (the Dyson orbital). The correct way to compare ionization probabilities of a group of MO in terms of their orbital energies is by looking at the ionization channels that can be approximated by the removal of an electron from the said MO where their ionization potentials are approximately calculated as an orbital energy of this MO.

### v3

@@ IV.4 N2, HCOOH, and CH3F
Figure 6: (a) and (b): $|G_{00}|^{2}$ of CH3F obtained by OE-WFAT using the two degenerate HOMO of the neutral molecule , respectively. (c) and (d): $|G_{00}|^{2}$ of CH3F obtained by ME-WFAT using the two degenerate ground states of the cation , respectively. These degenerate states are constructed at every orientation angle so that they diagonalize one-particle (in OE-WFAT) or all-particle (in ME-WFAT) dipole matrix.
@@
As the last example, we examine CH3F, a molecule having an equilibrium geometry that possesses a $C_{3v}$ symmetry, where the axis of the three-fold rotational symmetry coincides with the CF bond. The C–F and C–H bond lengths are $1.360$ and $1.080$, respectively, with the HCF angle of $108.93^{\circ}$. Its HOMO is doubly degenerate and can be identified based on their reflection symmetry with respect to one of the three FCH planes. Unlike the degenerate HOMO of OCS and O2, the two degenerate HOMO of CH3F do not diagonalize the dipole matrix, therefore, in view of the perturbation theory, a diagonalization of the dipole matrix in the subspace of degenerate HOMO (in OE-WFAT) or in the subspace of degenerate CH3F+ ground state (in ME-WFAT) must be performed at every orientation angle. The eigenvectors are then used in the WFAT integral formula—for ME-WFAT, these eigenvectors are used as $|\Psi^{+}\rangle$ wherever it appears. In practice, these dipole-diagonalizing states are never actually constructed at every orientation angle, which would otherwise require the expensive WFAT integrals (Eq. (III.3)) to be recomputed too. This is possible since structure factors are linear with respect to $|\Psi^{+}\rangle$. Therefore, the transformation can be applied to the structure factors of the non-diagonalizing states, avoiding the recomputation of WFAT integrals.
Fig. 6(a) and 6(b) show the $|G_{00}|^{2}$ structure factors obtained by OE-WFAT using the two degenerate HOMO of CH3F that diagonalize the one-particle dipole matrix. While Fig. 6(c) and 6(d) are the ME-WFAT counterparts obtained using the two degenerate ground states of CH3F+ that diagonalize the total (all-particle) dipole matrix. Both OE-WFAT and ME-WFAT structure factors in Fig. 6 exhibit the three-fold rotational symmetry of the molecule. We note in passing that the individual state’s structure factors calculated without diagonalizing the dipole matrix do not exhibit the three-fold rotational symmetry but still lead to identical total structure factor as those in Fig. 6.
Apart from the intensity, the most noticeable difference between the OE-WFAT and ME-WFAT results in Fig. 6 are the sizes and locations of the lobes, with ME-WFAT structure factors having smaller lobes than OE-WFAT and are located at a smaller $\beta$. Upon closer inspection, it can also be seen that the three lobes in the ME-WFAT results are more uniform in shape than those in OE-WFAT. We attribute this observation to the direction of the dipole moment vectors. Due to the approximate nature of the electronic structure computation, the dipole vectors of the degenerate CH3F HOMO (as used by OE-WFAT) are numerically less parallel to the $C_{3}$ axis of the molecule than the dipole vectors of the neutral’s ground state and the cation’s degenerate ground states (as used by ME-WFAT).
@@ Appendix A INTEGRAL FORMULAE FOR SINGLE-DETERMINANT WAVE FUNCTIONS
We note that the $\mathcal{P}(i)$’s encode the contribution of orbitals to the ionization rate through the Dyson orbital (see Eqs. (63) and (59)). On the other hand, orbital energy is also often used as a measure of contribution of that orbital to the rate. Nevertheless, orbital energy and $\mathcal{P}(i)$ are independent of each other. But they independently determine the overall contribution of an orbital to the rate—the energies affect the asymptotic behavior through the exponential tail, while the $\mathcal{P}(i)$’s determine the shape of the ionizing orbital, that is, the Dyson orbital.